Exercise 5-58 (LO5-4)

Miguel Cabrera of the Detroit Tigers had the highest batting average in the 2015 Major League Baseball season. His average was 0.477. So assume the probability of getting a hit is 0.477 for each time he batted. In a particular game, assume he batted four times.

  1. This is an example of what type of probability?

  • Classical

  • Empirical

  • Joint

  1. What is the probability of getting four hits in a particular game? (Round your answer to 3 decimal places.)

  1. What is the probability of not getting any hits in a game? (Round your answer to 3 decimal places.)

  1. What is the probability of getting at least one hit? (Round your answer to 3 decimal places.)
Answer #1

Ans:

a)Subjective

(Subjective probability is a type of probability derived from an individual's personal judgment about whether a specific outcome is likely to occur. It contains no formal calculations and only reflects the subject's opinions and past experience.)

b)

P(x=4)=0.477^4=0.052

c)

P(x=0)=(1-0.477)^4=0.075

d)

P(at least one)=1-0.075=0.925